# FlowHFT: Flow Policy Induced Optimal High-Frequency Trading under Diverse Market Conditions

### v1

@@ 4.3 Generation of State-Action Pairs
We generate state-action pairs for each market generated. For each parameter combination, we simulate $100$ episodes. Each episode consists of $1/dt$ time steps (e.g., $100$ steps if $dt=0.01$), each timestamp yields a state-action pair from the expert strategy, here we collect the corresponding state-action pairs from the dominant stochastic strategy among our candidates(AS, GLFT, GLFT-Drift, PPO). In total, we have gathered 1.62 million state-action pairs. These generated state-action pairs serve as the training dataset for the diffusion policy, providing a broad coverage of market conditions and ensuring robust policy learning. This dataset serves as the training data for our diffusion policy framework.
5 Experiments
To evaluate FlowHFT’s adaptability and performance, our experiments were conducted under out-of-sample market conditions. This was achieved by systematically controlling market parameters in the test environments to ensure they differed from those encountered during the training phase. The research questions are: (1) Can FlowHFT effectively generalize strategies learned from expert demonstrations to novel market conditions? (2) Does the integrated fine-tuning mechanism demonstrably improve the performance of actions initially proposed by the pre-trained model? (3) Can the FlowHFT framework achieve improved profitability in these challenging out-of-sample situations?
5.1 Test Market Environment Setting
We set the underlying price process’s memory characteristics by setting the Hurst exponent (H) to values of 0.2, 0.5, and 0.8. A Hurst exponent of H=0.2 generates strongly mean-reverting market conditions, where price changes tend to rapidly reverse direction. Setting H=0.5 simulates a random walk process typical of efficient markets with no memory effect, serving as a crucial baseline. On the other hand, H=0.8 creates strongly persistent or trending markets, where price movements are more likely to continue in their current direction. Second, we introduce variations in the overall market directionality by specifying the drift parameter ($\mu$) as either 0, representing a market with no inherent long-term upward or downward trend, or 0.2, introducing a positive drift to simulate a generally bullish market sentiment. These parameter choices result in 2×3=6 distinct core market scenarios based on trend and memory properties. Within each of these six scenarios, we further differentiate the market microstructure by creating four specific situations based on combinations of market volatility (Vol) and order arrival rate (AR): High Vol/High AR (HH), High Vol/Low AR (HL), Low Vol/High AR (LH), and Low Vol/Low AR (LL). This subdivision is designed to capture a range of typical market states: ’HH’ represents active, potentially news-driven volatile markets; ’HL’ mimics risky, potentially gappy conditions with low liquidity; ’LH’ reflects stable, liquid markets ideal for high-frequency trading; and ’LL’ simulates quiet, perhaps stagnant market periods. This comprehensive matrix of 6×4=24 scenarios provides a rigorous testbed for evaluating the adaptability and performance of HFT strategies across diverse and representative market conditions.
To implement the fine-tuning component, we determine the necessary adjustment parameter using a specific portion of the test market data. Each test period is divided into a validation set and a subsequent trading phase. The validation set period is utilized to calibrate the fine-tuning parameter. The calculated parameter is then used during the trading phase to adjust or modulate the actions proposed by the original pre-trained flow matching model.
@@ 5.2 Evaluation Metrics
  • Profit and Loss (PnL): Measures the total profits generated by the trading strategy.
  • Sharpe Ratio (SR): Assesses the risk-adjusted return.
  • Maximum Drawdown (MDD): Evaluates the risk by measuring the largest peak-to-trough decline in PnL.
@@ 5.3 Results Analysis
Looking at the results presented from Table 1 to Table 6, a consistent pattern emerges across the various simulated scenarios. Firstly, the GLFT model typically yields superior results compared to the AS model. Critically, our pre-trained flow matching policy model demonstrates performance levels comparable to the GLFT instructor. This finding is significant as it indicates that the pre-trained model successfully learned and internalized effective strategies from its instructors, even within these challenging stochastic environments.
Examining the performance results corresponding to the fine-tuned Flow Matching policy, a clear and significant trend emerges. This adapted model consistently and significantly outperforms both the traditional baseline models (AS and GLFT) and the initial pre-trained model across the tested environments. This finding is paramount, as it validates the second core objective of our framework: we not only demonstrate the ability to effectively learn from expert instructors but also showcase the capacity to surpass their performance. This success is attributed to the fine-tuning process, which empowers the model to "sense" the prevailing market conditions via the validation set calibration and subsequently adjust its behavior, specializing its strategy for the specific market environment encountered.
Table 1: Parameters for Market Conditions: Hurst Exponent $H=0.5$, Drift Rate $\mu=0$, volatility ($\sigma$: {0.02, 0.25}) and arrival rate ($\mu$: {25, 50}). The linear finetuning (ax+b) are the following: HH: action+[0.15,0.24], HL: action+[0.1,0.2], LH: action+[-0.1,0.1], LL: action+[-0.3,0.3],, these parameters can be chosen from an interval, just pick a point as representation The evaluation metrics are tested out of sample, with each methods are tested 1 million trails to obtain significance and convergence.
@@
WhiteNoise 1.99 0.06 28.90 0.99 0.04 19.51 2.10 0.28 2.87 1.09 0.19 1.98
AS 24.16 0.08 245.49 13.51 0.09 127.70 25.18 0.92 8.02 13.61 0.63 6.98
GLFT 25.11 0.36 61.39 13.55 0.23 53.26 25.21 0.99 7.45 13.62 0.66 6.59
Tables 4-7 expand the volatility and arrival rate options to investigate a broader spectrum of unseen scenarios. (Volatility, $\sigma$: {0.05, 0.2}) and arrival rates, $\mu$: {30, 60})
WhiteNoise 2.41 0.07 56.73 1.27 0.05 39.31 2.49 0.24 14.23 1.31 0.17 9.94
AS 28.82 0.14 346.8 16.23 0.13 216.69 29.28 0.78 48.86 16.16 0.54 40.22
Finetune Flow Policy 29.72 0.36 136.52 15.80 0.23 114.62 31.43 1.06 36.25 16.82 0.75 29.33
@@
WhiteNoise 2.48 0.10 11.72 1.31 0.07 8.13 2.51 0.28 3.27 1.32 0.20 2.32
AS 27.89 0.17 71.05 15.88 0.16 44.13 29.17 0.91 8.29 16.10 0.63 7.55
GLFT 29.00 0.57 21.05 15.77 0.36 19.88 29.21 1.10 6.63 16.07 0.73 6.38
WhiteNoise 2.48 0.10 11.92 1.32 0.07 8.27 2.51 0.26 3.42 1.33 0.18 2.42
AS 27.83 0.17 72.41 15.86 0.15 44.95 29.15 0.83 8.70 16.09 0.58 7.92
GLFT 28.98 0.56 21.46 15.76 0.35 20.23 29.19 1.00 6.95 16.07 0.67 6.68
@@ 6 Conclusion
In conclusion, we propose a novel imitation learning framrwork FlowHFT, the first application of flow matching policy in a financial stochastic control task. We have confirmed its capability to learn effective trading strategies in stochastic environments. A key strength of FlowHFT is its ability to integrate knowledge from expert strategies across diverse market scenarios into a single model, enabling prompt adaptation to prevailing conditions. Furthermore, the incorporated grid search mechanism successfully refines actions, especially in situations where those experts were suboptimal. FlowHFT thus represents a promising advancement in developing adaptive and high-performing HFT strategies.

### v2

@@ 4.2 Generation of State-Action Pairs
To generate diverse expert demonstrations for FlowHFT, we utilized a set of expert candidates, including: (1) the Avellaneda-Stoikov (AS) model avellaneda2008high , a foundational stochastic control framework for optimal quoting that considers inventory risk and market volatility; (2) the Guéant-Lehalle-Fernandez-Tapia (GLFT) model gueant2013dealing , which extends such optimal control approaches, often incorporating aspects such as order flow dynamics; (3) a modified GLFT model incorporating a price drift component, designed to adapt strategies in markets exhibiting directional trends; and (4)a model-free Reinforcement Learning (RL) agent trained with Proximal Policy Optimization (PPO) schulman2017proximal , which learns a policy through direct interaction or simulation to maximize a specified reward function. The detailed mathematical formulations, assumptions, and parameterization specifics for each of these expert models are provided in Appendix A.
@@
We generate state-action pairs for each market generated. For each parameter combination, we simulate $100$ episodes. Each episode consists of $1/dt$ time steps (e.g., $100$ steps if $dt=0.01$), each timestamp yields a state-action pair from the expert strategy, here we collect the corresponding state-action pairs from the dominant stochastic strategy among our candidates(AS, GLFT, GLFT-Drift, PPO). In total, we have gathered 3.24 million state-action pairs. These generated state-action pairs serve as the training dataset for the diffusion policy, providing a broad coverage of market conditions and ensuring robust policy learning. This dataset serves as the training data for our diffusion policy framework.
5 Experiments and Results Analysis
Our investigation is designed to critically evaluate FlowHFT’s advantages over existing approaches and its capabilities to earn profit. Specifically, we address the following research questions: RQ1: Can FlowHFT effectively generalize strategies learned from expert demonstrations to new, unseen market conditions? RQ2: Does the fine-tuning mechanism improve the performance of actions initially proposed by the pre-trained model? RQ3: Can the FlowHFT framework achieve greater profitability than traditional strategies?
5.1 Out-of-sample Market Environment Settings
To evaluate FlowHFT’s performance across a spectrum of out-of-sample market environments, we systematically configured distinct test conditions by setting key market parameters to differ from those used during training (see appendix C ). Firstly, to simulate varied stock price trends, we controlled the Hurst exponent, which governs the long-term memory and persistence patterns of the price series, and the drift term, which dictates its underlying direction. Secondly, based on these trend characteristics, we further differentiated the market microstructure by creating four specific situations based on combinations of market volatility (Vol) and overall order arrival rate (AR): (1) High Vol/High AR (HH), representing active, potentially news-driven volatile markets; (2) High Vol/Low AR (HL), mimicking risky market; (3) Low Vol/High AR (LH), reflecting stable, liquid markets; and (4) Low Vol/Low AR (LL), simulating quiet market periods. For each of these market situations, we perform traditional strategies as well as FlowHFT to compare their performance.
@@ 5.2 Evaluation Metrics
  • Profit and Loss (PnL): Measure the total change in the value over a specific period, reflecting the aggregate percentage gain or loss.
  • Sharpe Ratio (SR): It helps investors understand how much excess return an investment generated for each unit of risk it undertook. A higher Sharpe Ratio generally indicates a better risk-adjusted performance.
  • Maximum Drawdown (MDD): Quantify the largest percentage decline in the value of a strategy from a previous peak to a subsequent trough
The detailed calculations of these metrics are provided in Appendix B
@@ 5.3 Results Analysis
Looking at the results presented from Table 1 to Table 6, a consistent pattern emerges across the various market scenarios. Firstly, the GLFT model typically yields superior results compared to the AS model. Critically, our pre-trained flow matching policy model demonstrates performance levels comparable to the GLFT instructor. This finding is significant to address the RQ1. It indicates that the pre-trained model successfully learned and internalized effective strategies from its instructors, even within these challenging stochastic environments.
In response to RQ2, we examine the results corresponding to the fine-tuned Flow Matching policy. This adapted model typically and significantly outperforms both the traditional benchmark models and the initial pre-trained model across the tested environments. This finding validates the second core objective of our framework: we not only demonstrate the ability to effectively learn from expert instructors but also showcase the capacity to surpass their performance. This success is attributed to the fine-tuning process, which empowers the model to "sense" the prevailing market conditions via the validation set calibration and subsequently adjust its behavior, specializing its strategy for the specific market environment encountered.
Note that our experimental design includes highly volatile market environments, which are characterized by the potential for sudden and significant price movements. We show that FlowHFT exhibits robustness in these challenging conditions, particularly when compared with traditional RL methods. This indicates that generating sequences of actions rather than purely single-step decisions can effectively incorporate a longer-term planning perspective. This approach allows the model to better anticipate downstream consequences, make more considered adjustments over multiple steps, and mitigate the risk of compounding errors.
In response to RQ3, we take analyses of the performance metrics to obtain significant insights into the profit-generating capabilities across different market conditions. The Finetuned Flow Policy consistently demonstrates strong PnL (Profit and Loss) figures, notably achieving the highest PnL of 26.79 in the "Low Volatility & High Demand" scenario and 25.82 in the "High Volatility & High Demand" scenario. This indicates its robust ability to capitalize on profit opportunities in both highly active and more stable, liquid markets. Compared to the traditional models, the Finetuned Flow Policy generally matches or surpasses the PnL of AS, GLFT, and GLFT-drift in most scenarios, particularly excelling in high demand situations. The RL-PPO model, while adaptive, consistently shows lower PnL figures across all tested environments, underscoring the superior profit extraction of the FlowHFT framework.
Table 1: Parameters for Market Conditions: Hurst Exponent $H=0.5$, Drift Rate $\mu=0$, volatility ($\sigma$: {0.02, 0.25}) and arrival rate ($\mu$: {25, 50}). The linear finetuning (ax+b) are the following: HH: action+[0.15,0.24], HL: action+[0.1,0.2], LH: action+[-0.1,0.1], LL: action+[-0.3,0.3], these parameters can be chosen from an interval, just pick a point as representation The evaluation metrics are tested out of sample, with each methods are tested 1 million trails to obtain significance and convergence.
@@
The FlowHFT framework, with its flow matching policy, is designed to synthesize knowledge from numerous expert instructors. While our extensive experiments confirm its learning capabilities across many experts, we conduct a analysis to demonstrate its scaling law: that is, how performance improves as the model learns from an increasing number of distinct expert strategies. The results of this scaling analysis are presented in Figure 5.
Figure 5: Scaling analysis showing the relationship between the Number of Instructors (expert models) used in training FlowHFT and its resulting performance, as measured by Sharpe Ratio (blue bars, left y-axis) and Maximum Drawdown (red line, right y-axis).
Figure 5 illustrates the scaling properties of our FlowHFT framework, supporting the hypothesis that the model benefits from exposure to an increasing number of expert strategies. Two key observations underscore this scaling law. First, with an increase in the number of expert instructors used during training, the framework’s performance consistently improves over configurations with fewer experts. This progressive improvement indicates that FlowHFT effectively learns and integrates new, valuable strategic information from additional experts. Second, this enhanced learning translates into more robust risk-adjusted returns; not only does the Sharpe Ratio increase, but the Maximum Drawdown (red line) simultaneously shows a general decline. This continuous improvement in both return and risk metrics as more expert knowledge is incorporated suggests that FlowHFT is adept at leveraging a broader set of specialized strategies to find more consistently superior actions across diverse market conditions.
@@ 6 Conclusion
In conclusion, we propose a novel imitation learning framework FlowHFT, the first flow matching policy-based approach in a financial stochastic control task. Our findings confirm that it has ability to learn effective trading strategies in various stochastic environments in one model, thus enabling prompt adaptation to prevailing market conditions. Furthermore, the incorporated grid search mechanism successfully refines actions, especially in situations where these experts were suboptimal. In terms of profit generation, FlowHFT’s advantage over traditional models is evident through its achievement of greater cumulative returns, a superior Sharpe ratio indicating better risk-adjusted performance, and a reduced maximum drawdown signifying lower downside risk. Therefore, FlowHFT represents a promising advancement in the development of adaptive and high-performance HFT strategies.
@@ A.4 Proximal Policy Optimization (PPO) for High-Frequency Market Making
Appendix B Calculation of Evaluation Metrics
1. Cumulative Return (CR)
The Cumulative Return measures the total percentage change in the value of an investment over a specific period.
If $r_{i}$ is the return for period $i$, and there are $n$ periods:
$\text{CR}=\left(\prod_{i=1}^{n}(1+r_{i})-1\right)\times 100\%$
where:
  • $r_{i}$: Return for period $i$.
  • $n$: Total number of periods.
2. Sharpe Ratio (SR)
The Sharpe Ratio measures the risk-adjusted return of an investment or strategy. It quantifies the average return earned in excess of the risk-free rate per unit of volatility or total risk.
$\text{SR}=\frac{\mathbb{E}[R_{p}-R_{f}]}{\sigma_{p}}$
where:
  • $\mathbb{E}[R_{p}-R_{f}]$: The expected value of the excess return of the portfolio over the risk-free rate.
  • $\bar{R}_{p}$: The average (sample mean) return of the portfolio or investment over a period.
  • $\bar{R}_{f}$: The average (sample mean) risk-free rate of return over the same period (e.g., yield on short-term government bonds).
  • $\sigma_{p}$: The standard deviation of the portfolio’s excess returns ($R_{p}-R_{f}$). (Sometimes, the standard deviation of the portfolio’s returns, $\sigma(R_{p})$, is used as an approximation, especially if $R_{f}$ is constant or has very low volatility).
3. Maximum Drawdown (MDD)
The Maximum Drawdown is the largest percentage drop in the value of an investment or portfolio from a previous peak to a subsequent trough over a specific period. Let $V(t)$ be the value of the portfolio at time $t$ over the period $[0,T]$.
First, define the running peak value up to time $t$ as $M(t)$:
$M(t)=\max_{0\leq\tau\leq t}V(\tau)$
Then, the drawdown $D(t)$ at any time $t$ is:
$D(t)=\frac{M(t)-V(t)}{M(t)}$
The Maximum Drawdown (MDD) is the maximum value of $D(t)$ over the entire period:
$\text{MDD}=\max_{0\leq t\leq T}D(t)$
This MDD is usually expressed as a positive percentage (e.g., an MDD of 0.25 means a 25% maximum loss from a peak).
where:
  • $V(t)$: Value of the investment at time $t$.
  • $T$: The end of the observation period.
  • $\tau$: A time index variable ranging from 00 to $t$.
Appendix C Parameters for Test Environemnt
Figure 6: Parameters settings of testing market environement
Appendix D Results for Different Market Conditions
AS 24.16 0.08 245.49 13.51 0.09 127.70 25.18 0.92 8.02 13.61 0.63 6.98
GLFT 25.11 0.36 61.39 13.55 0.23 53.26 25.21 0.99 7.45 13.62 0.66 6.59
Tables 4 to 6 expand the volatility and arrival rate options to investigate a broader spectrum of unseen scenarios. (Volatility, $\sigma$: {0.05, 0.2}) and arrival rates, $\mu$: {30, 60})
AS 28.82 0.14 346.8 16.23 0.13 216.69 29.28 0.78 48.86 16.16 0.54 40.22
Finetuned Flow Policy 29.72 0.36 136.52 15.80 0.23 114.62 31.43 1.06 36.25 16.82 0.75 29.33
GLFT 29.00 0.57 21.05 15.77 0.36 19.88 29.21 1.10 6.63 16.07 0.73 6.38
GLFT 28.98 0.56 21.46 15.76 0.35 20.23 29.19 1.00 6.95 16.07 0.67 6.68
Appendix E Limitations
While FlowHFT demonstrates promising capabilities, certain limitations in the current study should be acknowledged. One limitation of the current study is its focused scope, with the FlowHFT framework being developed and evaluated exclusively for a single financial asset trading on a single exchange. While this controlled setting allows for a detailed analysis of the framework’s core mechanics and learning capabilities, we can further address the broader complexities and rich dynamics present in multi-asset or multi-exchange HFT operations. For instance, this study can explore how FlowHFT would manage inter-asset correlations, portfolio-level risk diversification, or the strategic allocation of capital across different instruments. Similarly, the challenges and opportunities arising from operating across multiple exchanges—such as navigating varying fee structures, latency differences, diverse order book microstructures, and potential inter-exchange arbitrage—remain outside the current investigation. Therefore, a significant and crucial avenue for future research is the extension and evaluation of the FlowHFT framework to multi-asset and multi-exchange environments. This would involve adapting the state and action representations, potentially incorporating mechanisms to learn cross-asset dependencies, and developing policies capable of sophisticated order routing and execution across different trading venues. Successfully addressing these aspects would substantially broaden FlowHFT’s applicability and provide a more comprehensive validation of its potential as a robust solution for real-world high-frequency trading.